Dear Sir/Madam,
I cannot figure out these simultaneous equations. Please help and show all working out.
(12-p)/p=12/b
12/p=(24-b)/b

Answer by malglu(63)

these are proportions so you can multiply the denominators on each side by the numeratpr of the other, to get rid of the fracyions
you get
12b -bp =12p
and
12b = 24p -pb
rearrange these so that you have the b term on one side and the other two terms on the other. then minus one from the other, to get
24b/36 = p
i.e. p = 2/3rds b
if you substuitue this into one of the equations above, for eveyr time you have a 'p' then you will now get an equation which only has b's in it.
2b^2 -12b = 0
this can now be solved
b=0 or b =6